OpenStudy (anonymous):

how do u find the least common multiple of 22 and 46

OpenStudy (cwrw238):

22 = 2 * 11 46 = 2 * 23 lcm = 2*11*23 = 506 2 (is included only once in calculation because it is common to both factors )

OpenStudy (anonymous):

but theyy both need a lcm and for both of them to do tht they have to have a lcm tht is the same

OpenStudy (cwrw238):

an lcm is a multiple of 2 or more numbers

OpenStudy (anonymous):

ik tht but wat is it then

OpenStudy (cwrw238):

sorry i dont follow you - do you want the definition of lcm - least common multiple? lcm of 22 and 46 = 506

OpenStudy (anonymous):

k thnx bye

OpenStudy (cwrw238):

bye